OpenStudy (anonymous):

For tax reasons, I need to create a rectangular vegetable patch with an area of exactly 50 sq. ft. The fencing for the east and west sides costs $4 per foot, and the fencing for the north and south sides costs only $2 per foot. What are the dimensions of the vegetable patch with the least expensive fence?

OpenStudy (radar):

Let x = length of north or south side. 50/x = length of east or west side. 2x($2) = cost of north or south side. 2(50/x)($4) = cost of east/west side. total cost =4x + 400/x Cost = C C=4x+400/x Take the derivative C'=4-400/x^2 For minima set to = 0 4-400/x^2 = 0 Multiply thru by X^2 4x^2 - 400 = 0 Divide thru by 4 x^2-100=0 Factor (difference of two perfect squares) (x+10)(x-10) = 0 x=10 disregarding negative value. North and South side is 10 ft ea. 50/x=50/10=5 East/West side is 5 ft. each.

OpenStudy (anonymous):

Thank you so much I really appreciate it. Thank you brother.

OpenStudy (radar):

ur welcome, good luck.